OpenStudy (anonymous):

What point is the tangent line to the curve y =x^3 +5 perpendicular to the line x+3y=2

OpenStudy (anonymous):

I got the slope of the normal line and points of tangency. I don't know what I need to do next..

OpenStudy (anonymous):

I'm not really sure with my points of tangency tho

OpenStudy (vivek3461):

Slope of the line x + 3y = 2 is -1/3 So the slope of the perpendicular line is 3 ( product slopes should be -1) This is required slope of the tangent line From the given curve equation, y = x^3 + 5, Slope of the tangent line for this curve is dy / dx = 3.x^2 This slope should be 3, So 3.x^2 = 3 x^2 = 1 x = +1,-1 Substituting these values of x in the curve equation we get, the two points (1,6) and (-1,4)

OpenStudy (anonymous):

ooh so we can equate the slope of the tangent line of the curve and the slope of the normal line?

OpenStudy (anonymous):

i mean perpendicular line?

OpenStudy (vivek3461):

Yes you are right

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

I have another question related to this.

OpenStudy (anonymous):

why do we have to substitute our point of tangency to the curve and not on the equation of the line?